Suppose that you have 10 green cards and 5 yellow cards. The cards are well shuffled. You randomly draw two cards with replacement. Round your answers to four decimal places.
G1 = the first card drawn is green
G2 = the second card drawn is green
a. P(G1 and G2) =
b. P(At least 1 green) =
c. P(G2|G1) =

d. Are G1 and G2 independent?

Answer by rothauserc(4718)

there are a total of 15 cards, 10 green, 5 yellow
:
a. P(G1 and G2) = 10/15 * 10/15 = 2/3 * 2/3 = 4/9 approximately 0.4444
:
b. P(at least 1 is green) = (2/3 * 1/3) + (2/3 * 2/3) = 6/9 = 2/3 approximately 0.6667
:
c. P(G2|G1) = 2/3 approximately 0.6667
:
independent because you are replacing a card after it is drawn
: